I am trying to understand how to divide fractions.
can you help me with this problem?
2/4 divided by 1/5=
thank you

Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(69443)

2/4 divided by 1/5
-------------
Invert the divisor (1/5) and multiply
= (2/4)*(5/1)
= 10/4
= 5/2

Answer by richwmiller(17219)

let's look at some examples first and see what we notice.
if we divide 20 by 5 we get 4
20/5=4 but 20 *1/5 also equals 4
so dividing by a number is the same as if we multiply by the reciprocal
so if we divide by 1/5 we can also multiply by 5
if we divide by 5 we can multiply by 1/5
2/4 divided by 1/5 is the same as multiplying by 5
2/4*5=10/4
BTW 2/4 -1/2
so 1/2*5=10/4=2.5 or 2 1/2